OpenStudy (anonymous):

Multiply and simplify. 2^8 * 2^9 type your answer using exponential notation

jimthompson5910 (jim_thompson5910):

Add the exponents, so 2^8 * 2^9 = 2^(8+9) = ???

OpenStudy (anonymous):

That where I was stuck :) so is 2^17?

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

Thank you I was going to multiply.

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

2^8+9=2^17=512

jimthompson5910 (jim_thompson5910):

it's not 512, they don't want the answer in that form anyway

OpenStudy (anonymous):

the answer has to be in exponential notation BUt thank you :)

OpenStudy (anonymous):

ok, here is 2^8+9=2^17